# Uncertainty Quantification on Graph Learning: A Survey

### v2

@@ Abstract
Graphical models have demonstrated their exceptional capabilities across numerous applications, such as social networks, citation networks, and online recommendation systems. Despite these successes, their performance, confidence, and trustworthiness are often limited by the inherent randomness of data in nature and the challenges of accurately capturing and modeling real-world complexities. This has increased interest in developing uncertainty quantification (UQ) techniques tailored to graphical models. In this survey, we comprehensively examine these existing works on UQ in graphical models, focusing on key aspects such as foundational knowledge, sources, representation, handling, and measurement of uncertainty. This survey distinguishes itself from most existing UQ surveys by specifically concentrating on UQ in graphical models, particularly probabilistic graphical models (PGMs) and graph neural networks (GNNs). We elaborately categorize recent work into two primary areas: uncertainty representation and uncertainty handling. By offering a comprehensive overview of the current landscape, including both established methodologies and emerging trends, we aim to bridge gaps in understanding and highlight key challenges and opportunities in the field. Through in-depth discussion of existing works and promising directions for future research, we believe this survey serves as a valuable resource for researchers, inspiring them to cope with uncertainty issues in both academic research and real-world applications.
@@ I Introduction
Graphical models have emerged as fundamental tools in machine learning due to their capability to capture and manage complex relational dynamics among variables. These models have found extensive applications in various domains and tasks, such as spam detection in online review systems [1, 2], relationship extraction in social networks [3, 4, 5, 6], interest exploration in recommendation systems [7, 8, 9, 10], and drug discovery in protein-protein interaction networks [11, 12, 13]. Unlike traditional models, where variables are assumed to be independently and identically distributed (i.i.d.), graphical models reveal intricate relationships and conditional dependencies. Considering both individual variables and their interconnections, graphical models improve predictive performance in complex data environments and tasks.
@@
However, uncertainties are ubiquitous and can arise from various sources, including inherent data randomness, model training errors, and unexpected test data distributions [14]. These uncertainties hinder the application of graphical models in critical fields where precise risk assessment and decision-making under uncertainty are essential. Quantifying uncertainty (UQ) provides confidence measures and explanations alongside predictions, enhancing the reliability and safety of models in practice. Thus, many works have focused on developing UQ methods specifically for graphical models, allowing their applications in domains such as healthcare [15] and autonoumous driving [16].
@@
This survey explores various types of uncertainties in two primary types of graphical models: graph neural networks (GNNs) and probabilistic graphical models (PGMs). GNNs extend the successes of deep learning to graph-structured data but face challenges such as noisy links, missing nodes, or mislabeled data. Recent developments have incorporated uncertainty quantification into the architectures and training algorithms of GNNs, improving their robustness and interoperability. PGMs, including Bayesian Networks and Markov Random Fields, utilize more rigorous probabilistic formulations to effectively represent uncertainty and are particularly effective in scenarios requiring joint probability modeling and causal relationship analysis.
Figure 1: Overview of the uncertainty quantification: the survey introduces the source of uncertainty in Sec. III. Then, the state-of-the-art methods proposed for uncertainty representation and handling are demonstrated in Sec. IV and Sec. V, respectively. To assess the results of UQ, the evaluation protocols on UQ will be presented in Sec. VI.
Existing surveys on UQ can be categorized based on their focus, such as their specific application domain (e.g., UQ for graph learning), and the breadth of their coverage, ranging from comprehensive overviews of UQ methods to in-depth explorations of particular UQ techniques. The survey most closely related to the present survey is [14], which also focuses on graph learning and UQ. Similarly to our survey, [14] investigates the sources of uncertainty, evaluation methods, and various techniques for handling uncertainty in graph learning. However, a key distinction between [14] and our survey is that the former does not include a discussion on the representation methods for uncertainty. In broader application domains, some surveys provide comprehensive coverage of UQ methods. For example, [17, 18, 19, 20] explore uncertainty in deep learning, while [21] examines uncertainty in scientific machine learning. These surveys provide valuable insights into UQ across different application domains. However, they do not specifically delve deep into the uncertainty in graph learning. Furthermore, several surveys offer an in-depth analysis of specific aspects of UQ, such as sources of uncertainty [22, 23], optimization under uncertainty [24], noise in machine learning [25], but do not focus on uncertainty in graph learning. Despite the valuable contributions of the existing work mentioned above, they collectively lack a comprehensive overview of uncertainty quantification on graph learning.
This survey aims to review various strategies and methodologies to model, measure, and mitigate uncertainty in GNNs and PGMs. By examining recent research, including state-of-the-art works, the survey provides a comprehensive understanding of how current research addresses uncertainty and forecasts future enhancements to increase their efficacy. Specifically, this survey organizes the literature on uncertainty in graphs into four main parts, as demonstrated in Fig. 1. Specifically, we first provide the preliminary knowledge concerning PGMs and GNNs in Sec. II, and introduce the sources of uncertainty in Sec. III. We then discuss the methods used for uncertainty representation in Sec. IV and uncertainty handling in Sec. V. Table. II categorizes these UQ techniques by their properties. For instance, we are interested in whether they involve model training or post-hoc, whether they are based on uncertainty or deterministic parameters, and whether they have impacts on accuracy performance. Additionally, we will demonstrate widely used metrics for uncertainty evaluation in Sec. VI. Finally, the conclusion of the survey will be made in Sec. VIII.
@@ II Preliminaries
$Y_{i}\in\mathcal{Y}$ Ground truth of node $V_{i}$.
$\hat{Y}_{i}\in\mathcal{Y}$ Approximation of node $V_{i}$.
@@ II-A Basic Notations for Graphs
We summarize the formal notations for key definitions and variables in Table I. Consider a graph $G=(\mathcal{V},\mathcal{E})$ consisting of a set of $n$ nodes $\mathcal{V}=\{V_{1},\dots,V_{n}\}$, each representing a random variable, and each variable $V_{i}\in\mathcal{V}$ is associated with a feature vector $X_{i}$. The “feature” could be a prior distribution indicating the likelihood of belonging to class $k$ (typically in MRF cases), or a vector with practical meanings, such as bag-of-words (typically in GNN cases). Each edge in the set of edges $\mathcal{E}$ is a relationship between two variables, The set $\mathcal{N}(V_{i})$ collects direct neighbors of $V_{i}$, e.g., $\mathcal{N}(V_{i})=\{V_{j}:(V_{i},V_{j})\in\mathcal{E}\}$. In the node classification task, for each unlabeled variable $V_{i}$, the model aims to estimate its probability distribution $\hat{Y}_{i}$ and predict its label $k_{i}$ out of $K$ possible classes.
@@ II-B Probabilistic Graphical Model
A probabilistic graphical model (PGM) [26, 27] is widely used to represent the joint probability distribution of random variables, whose relations can be depicted in a graph. There are two distinct categories of PGMs, depending on whether the graph’s edges are directed or undirected. Specifically, a Bayesian Network is employed for modeling a directed graph, where edges represent causality. Conversely, a Markov Random Field (MRF) is utilized to model an undirected graph, where edges signify the correlation between nodes. An MRF factorizes the joint distribution $P(\mathcal{V})$ as
@@
where $Z$ normalizes the product to a probability distribution. $\phi(X_{i}=x_{i})$ denotes the prior probability of $X_{i}$ taking value $x_{i}$, independent of other variables. The compatibility $\psi(X_{i}=x_{i},X_{j}=x_{j})$ represents the preference of the pair $(x_{i},x_{j})$ occurring jointly, capturing the dependencies between the two variables. The marginal distribution $b(X_{i})$ for $V_{i}\in\mathcal{V}$ is
$b(X_{i}=x_{i})=\sum_{\mathcal{V}\setminus X_{i}}P(\mathcal{V}\setminus X_{i},X_{i}=x_{i}).$ (2)
@@ II-C Graph Neural Networks
Unlike traditional neural networks that operate on grid-like data (such as tabular vectors, images, or sequences), graph neural network (GNN) could learn from complex relationships and dependencies in graphs. Formally, for the $l$-th layer of an $L$-layers GNN, the node representation $\mathbf{h}_{i}^{(l)}$ for node $V_{i}$ is calculated by
$\mathbf{a}^{(l)}_{i}$ $=$ $\textnormal{AGG}\left(\left\{\mathbf{h}^{(l-1)}_{j}|V_{j}\in\mathcal{N}(V_{i})\right\}\right),$ (3)
$\mathbf{h}^{(l)}_{i}$ $=$ $\textnormal{UPDATE}\left(\mathbf{a}^{(l)}_{i},\mathbf{h}^{(l-1)}_{i}\right).$ (4)
The AGG function aggregates the messages sent from all neighbors $V_{j}\in\mathcal{N}(V_{i})$ to $V_{i}$. It could operate the element-wise average, maximum over the message set, or employ a parametric layer for aggregation [28]. The UPDATE function combines representations of node $X_{i}$ and its neighbors. $\mathbf{h}^{(0)}_{i}$ is the input node feature for node $X_{i}$. $\mathbf{h}_{i}^{(L)}$ can be softmaxed to a probability distribution for node-classification tasks, or be mapped to graph representation by a READOUT function for graph-level tasks. The parameters of GNN, $\boldsymbol{\theta}=\{\theta^{(l)}$, $l=1,\dots,L$, are trained in an end-to-end fashion.
@@ III Sources of Uncertainty
The total uncertainty in machine learning can be decomposed into aleatoric uncertainty and epistemic uncertainty based on their sources [29]. Aleatoric uncertainty, which is irreducible with increasing amount of data, represents the randomness inherent to the data generation process [30]. However, epistemic uncertainty can be reduced by obtaining more knowledge or data to estimate the ground truth of predictions better [31, 23]. The techniques for the disentanglement of aleatoric and epistemic uncertainty are introduced by [32], and [33] formulate mathematical expressions of these two kinds of uncertainties in the context of GNNs.
The identification of aleatoric and epistemic uncertainty must be built upon the context of the selected model and available data [34]. For example, flipping a coin is usually regarded as a purely random event, involving aleatoric uncertainty. However, if we know the entire physics process of flipping a coin, including the flipping force, the coin’s mass, volume, initial position, etc, we can simulate the process and predict whether it would be head or tail. Based on the simulation, the randomness can be regarded as epistemic uncertainty and can be reduced or even eliminated.
Focusing on supervised machine learning, we further explore the possible sources of uncertainty and how they impact prediction reliability. Let $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$ denote input feature variable and the output label variable with realizations $x$ and $y$ respectively. The underlying data distribution is $P_{XY}$ and the data generation mechanism is $f$ such that $y=f(x)$.
@@ III-A Aleatoric Uncertainty
Aleatoric uncertainty is only about data noises, so we do not need to consider the influence of model selection and training issues and can investigate this problem under the view of statistics. According to [22], when $f$ is conditioned on $X=x$, aleatoric uncertainty is the conditional probability distribution of $Y$, and its variance, $Var_{Y|X}(Y|x)$, can be used to quantify:
$Y|x\sim P_{Y|X}(\cdot|X=x).$ (5)
Four sources of aleatoric uncertainty are briefly introduced below. We refer the readers to  [22] for a more comprehensive and detailed taxonomy.
Omitted Features: Omitted features, which are investigated in the context of causal models [35], confounder analysis [36], and model sensitivity analysis [37], are significant sources of aleatoric uncertainty. The omission can be caused by unobservability of data or overlooking during data collection. Let $Z\in\mathcal{Z}$ denote an unobserved variable that influences the value of $Y$, we can see the expectation of $Y|x$ is a weighted integration of $\mathbb{E}_{Y|X,Z}$ on $Z$:
@@
In [22], the authors also claim that $Var_{Y|X,Z}(Y|x,z)\leq Var_{Y|X}(Y|x)$ with at least one $z\in\mathcal{Z}$, indicating that including the non-negligible variables will reduce aleatoric uncertainty.
Feature Errors: We will likely lose information with improper measurements, like low-resolution cameras, low-quality audio recorders, and uncalibrated gauges. These errors can cause biases and increased variance in feature collection. The authers of [25] reviewed 79 works on identifying and handling data noise. Here we let $Z$ represent feature variables by faultless measurements and $X$ denote those by inaccurate measurements, assuming that $Y$ is independent of $X|Z$. According to [22], the change of aleatoric uncertainty caused by noisy feature values is illustrated as
$Var_{Y|X}(Y|x)=$ $\mathbb{E}_{Z|X}[Var_{Y|Z}(Y|Z)|x]$ (7)
$\;+Var_{Z|X}(\mathbb{E}_{Y|Z}[Y|Z]|x).$
The first term of the right-hand side is the expectation of $Var_{Y|Z}(Y|Z)$ conditioned on $X=x$, and the second term is an increment in aleatoric uncertainty.
[…]
$Y|x\sim P_{Y|X}(\cdot|x)=\int_{z\in\mathcal{Z}}P_{Y|X,Z}(\cdot|x,z)P_{Z|X}(z|x)dz.$ (8)
Missing Data: Missing data indicates some entries in the dataset are incomplete [39]. If $(x_{i},y_{i})$ is the $i$-th sample of the dataset, $M_{i}$ is its corresponding missing indicator. $M_{i}=0$ means the sample is complete, and $M_{i}=1$ otherwise. A bias factor is explained in  [22] as $\frac{P(M=0|y,x)}{P(M=0|x)}$ to express the influence from incompleteness by
III-B Epistemic Uncertainty
Epistemic uncertainty is introduced by model selection and training, rooted in the lack of data and knowledge. In  [23], epistemic uncertainty is decomposed into model and approximation uncertainty. Let $f\in\mathcal{F}$ be the ground truth mapping relationship from $X$ to $Y$, $h\in\mathcal{H}\subset\mathcal{F}$ be the best possible predictor we can get, and $\hat{h}\in\mathcal{H}$ be the induced predictor in practice.
Model Uncertainty measures the difference between $f$ and $h$. This is because we can only explore a limited function space $\mathcal{H}$, which may not include the true $f$. Therefore, there can be a distance between the optimally trained predictor and the real mapping function. Taking $l$ as the loss function during training, $f$ and $h$ are explicitly defined in [23] as
@@ III-B Epistemic Uncertainty
Approximation Uncertainty counts for the uncertainty due to the limited number of training samples. Assuming we have $N$ samples available , the definition of $\hat{h}$ is listed as
$\hat{h}=\arg\min_{h\in\mathcal{H}}\frac{1}{N}\sum_{i=1}^{N}l(y_{i},h(x_{i})).$ (12)
It is obvious that, unlike aleatoric uncertainty, model uncertainty and approximation uncertainty are reducible by enlarging the function space $\mathcal{H}$ and increasing sample size $N$. We recommend [40, 41, 42] for the latest progress in predicting, sampling, and quantifying epistemic uncertainty.

### v3

@@ Abstract
Graphical models have demonstrated their exceptional capabilities across numerous applications, such as social networks, citation networks, and online recommendation systems. However, their performance, confidence, and trustworthiness are often limited by the inherent randomness in data and the challenges of accurately modeling real-world complexities. There has been increased interest in developing uncertainty quantification (UQ) techniques tailored to graphical models. In this survey, we comprehensively examine existing works on UQ for graphical models, focusing on key aspects such as the sources, representation, handling, and evaluation of uncertainty. This survey distinguishes itself from most existing UQ surveys by specifically concentrating on UQ in graphical models, including probabilistic graphical models (PGMs) and graph neural networks (GNNs). After reviewing sources of uncertainty, we organize the work using two high-level dimensions: uncertainty representation and uncertainty handling. By offering a comprehensive overview of the current landscape, including both established methodologies and emerging trends, we aim to bridge gaps in understanding key challenges and opportunities in UQ for graphical models, hoping to inspire researchers working on graphical models or uncertainty quantification to make further advancements at the cross of the two fields.
@@ I Introduction
Graphical models have emerged as fundamental tools in machine learning due to their capability to capture and manage complex relational dynamics among variables. These models have found extensive applications, such as spam detection in online review systems [1, 2], relationship extraction in social networks [3, 4, 5, 6], user interest exploration in recommendation systems [7, 8, 9, 10], and drug discovery in protein-protein interaction networks [11, 12, 13]. Unlike traditional models, where variables are assumed to be independently and identically distributed (i.i.d.), graphical models could reveal intricate relationships and conditional dependencies. By considering both individual variables and their interconnections, graphical models improve predictive performance in complex data environments and tasks.
@@
However, uncertainties are ubiquitous and can arise from various sources, including inherent data randomness, model training errors, and unexpected test data distributions [14]. These uncertainties hinder the deployment of graphical models in high-stakes domains where precise risk assessment and decision-making under uncertainty are vital, including healthcare [15, 16] and autonomous driving [17, 18]. Uncertainty quantification (UQ) provides calibrated confidence measures alongside predictions, thereby improving the reliability of the models in practice [19, 20, 21]. Thus, many studies focus on developing UQ techniques for graphical models, enabling reliable applications in critical fields such as climate risk forecasting [22] and intelligent control systems [23].
@@
This survey explores various types of uncertainties in two primary types of graphical models: probabilistic graphical models (PGMs) and graph neural networks (GNNs). PGMs, including Bayesian Networks and Markov Random Fields, utilize more rigorous probabilistic formulations to effectively represent uncertainty and are particularly effective in scenarios requiring joint probability modeling and causal relationship analysis. GNNs extend the successes of deep learning to graph-structured data but face uncertainty challenges such as noisy links, missing nodes, or mislabeled data. Recent developments have incorporated uncertainty quantification into the architectures and/or training algorithms of GNNs, improving their robustness and interoperability.
Figure 1: We first show an example from a recommendation system, where the user of interest (in gray) links to other users and contents with various relationships. A model’s predictions for the user’s preference under uncertainty are shown below, with or without uncertainty handling methods. Then, as detailed in Sec. III, the source of uncertainty can be decomposed into aleatoric and epistemic uncertainties. Last, we demonstrate the state-of-the-art methods proposed for uncertainty representation and handling, respectively. Specifically, uncertainty representation methods (refer to Sec. IV) include graph variational autoencoder, Bayesian graph neural network, and OOD epistemic uncertainty representation methods. Besides, as discussed in Sec. V, uncertainty handling methods such as OOD handling techniques, conformalization, and calibration, are shown in the figure.
Existing surveys on UQ provide comprehensive coverage of UQ methods, and offer valuable insights into UQ across various application domains. These surveys can be categorized based on specific application domains (e.g., UQ for graph learning) and the extent of their coverage, ranging from broad overviews of UQ methods [24, 25, 26, 27] to focused studies on particular areas such as scientific machine learning [28], sources of uncertainty [29, 30], optimization under uncertainty [31], and noise in machine learning [32]. Other surveys focus on specific paradigms, like integrating differential equations (DEs) with GNNs, which includes some discussion of uncertainty within the DE context [33]. However, none of these focus on uncertainty quantification on graph learning.
The survey most closely related to ours is [14], which provides a broad overview of UQ in graph learning. However, Wang et al. [14] focus primarily on downstream tasks and applications related to GNN uncertainty, rather than on the theoretical foundations of UQ methods themselves. In contrast, our work delves into the mathematical underpinnings of each UQ approach, presenting the key formulas and core concepts that offer a principled perspective on uncertainty in graphs. Moreover, whereas Wang et al. categorize methods by “single models versus ensemble models” and “deterministic models versus models with random parameters,” these labels obscure the conceptual differences between approaches. We instead organize UQ methods into “Uncertainty representation” (how uncertainty is modeled) and “Uncertainty handling” (how uncertainty is utilized or mitigated), which more clearly reflects the roles of different techniques. Finally, our survey incorporates advanced UQ for graphs methodologies, such as frameworks based on stochastic differential equations, which are entirely absent from [14].
Fig. 1 shows an example of uncertainty on a recommendation system, and highlights three key components of the survey: source, representation, and handling of uncertainties. For the organization, we first provide the preliminary knowledge concerning PGMs and GNNs in Sec. II, and introduce the sources of uncertainty in Sec. III. We then discuss the existing studies for uncertainty representation and uncertainty handling in Sec. IV and Sec. V, respectively. Furthermore, we demonstrate widely used metrics for uncertainty evaluation in Sec. VI. Finally, future directions and the conclusion of the survey are discussed in Sec. VII and Sec. VIII, respectively.
@@ II Preliminaries
$Y_{i}\in\mathcal{Y}$ True label of node $V_{i}$.
$\hat{Y}_{i}\in\mathcal{Y}$ Predicted label of node $V_{i}$.
@@ II-A Basic Notations for Graphs
We summarize the key notations in Table I. Consider a graph $G=(\mathcal{V},\mathcal{E})$ consisting of a set of $n$ nodes $\mathcal{V}=\{V_{1},\dots,V_{n}\}$, each representing a random variable, and each variable $V_{i}\in\mathcal{V}$ is associated with a feature vector $X_{i}$. The “feature” could be a prior distribution indicating the likelihood of belonging to class $k$ (typically in MRF), or a vector with practical meanings, such as bag-of-words (typically in GNN). Each edge in the set of edges $\mathcal{E}$ is a relationship between two variables. The set $\mathcal{N}(V_{i})$ collects direct neighbors of $V_{i}$, e.g., $\mathcal{N}(V_{i})=\{V_{j}:(V_{i},V_{j})\in\mathcal{E}\}$.
@@ II-B Probabilistic Graphical Model
A probabilistic graphical model (PGM) [34, 35] is widely used to represent the joint probability distribution of random variables, whose relations can be depicted in a graph. There are two distinct categories of PGMs, depending on whether the graph’s edges are directed or undirected. Specifically, a Bayesian Network (BN) uses a directed graph, where edges represent causality. Conversely, a Markov Random Field (MRF) uses an undirected graph, where edges signify the correlation between nodes. Both BN and MRF factorizes the joint distribution $P(\mathcal{V})$ as follows [35]:
$P(\mathcal{V})=\frac{1}{Z}\prod_{s\in\mathcal{S}}\zeta_{s}(X_{s}),$ (1)
where each factor $\zeta_{s}(X_{s})$ is a nonnegative function defined on a subset $X_{s}\subseteq\{X_{1},X_{2},\dots,X_{n}\}$, and $Z$ is for normalization. A BN defines $\zeta_{s}$ as the node’s probability distribution conditioned on their parents and naturally $Z=1$. Formally,
$P(\mathcal{V})=\prod_{i}^{n}P(X_{i}|Pa(X_{i})),$ (2)
where $Pa(X_{i})$ specifies parent nodes of $X_{i}$. An MRF factorizes the joint distribution $P(\mathcal{V})$ as
@@
where $Z$ normalizes the product to a probability distribution. $\phi(X_{i}=x_{i})$ denotes the prior probability of $X_{i}$ taking value $x_{i}$, independent of other variables. The compatibility $\psi(X_{i}=x_{i},X_{j}=x_{j})$ represents the probability of the pair $(x_{i},x_{j})$ occurring jointly, capturing the dependencies between the two variables. The marginal distribution $b(X_{i})$ for $V_{i}\in\mathcal{V}$ is
$b(X_{i}=x_{i})=\sum_{\mathcal{V}\setminus\{X_{i}\}}P(\mathcal{V}\setminus\{X_{i}\},X_{i}=x_{i}).$ (4)
@@ II-C Graph Neural Networks
Unlike traditional neural networks that operate on grid-like data (such as tabular vectors, images, or sequences), graph neural network (GNN) could learn from complex relationships in graphs. Formally, for the $l$-th layer of an $L$-layers GNN, the representation $\mathbf{h}_{i}^{(l)}$ of node $V_{i}$ is
$\mathbf{a}^{(l)}_{i}$ $=$ $\textnormal{AGG}\left(\left\{\mathbf{h}^{(l-1)}_{j}|V_{j}\in\mathcal{N}(V_{i})\right\};\theta^{(l)}\right),$ (5)
$\mathbf{h}^{(l)}_{i}$ $=$ $\textnormal{UPDATE}\left(\mathbf{a}^{(l)}_{i},\mathbf{h}^{(l-1)}_{i};\theta^{(l)}\right).$ (6)
The AGG function aggregates the messages sent from all neighbors $V_{j}\in\mathcal{N}(V_{i})$ to $V_{i}$, and it could operate as an element-wise average, or maximum over the message set, or employ a parametric layer for aggregation [36]. The UPDATE function combines representations of node $X_{i}$ and its neighbors. $\mathbf{h}^{(0)}_{i}$ is the input node feature for node $X_{i}$. The representations from the last layer $\mathbf{h}_{i}^{(L)}$ can be softmaxed to a probability distribution for node-classification tasks [6], used in a scoring function for link prediction tasks [37], or mapped to a graph representation by a READOUT function for graph-level tasks. The parameters of GNN, $\boldsymbol{\theta}=\{\theta^{(l)}$: $l=1,\dots,L$, are trained in an end-to-end fashion.
@@ III Sources of Uncertainty
The total uncertainty in machine learning can be decomposed into aleatoric uncertainty (AU) and epistemic uncertainty (EU) based on their sources [38]. [^1] Recent studies have challenged the traditional binary classification of uncertainty into aleatoric and epistemic types, suggesting that this dichotomy may be overly simplistic [39, 40]. Aleatoric uncertainty, which is irreducible with increasing amount of data, represents the randomness inherent to the data generation process [41]. On the other hand, epistemic uncertainty can be reduced by obtaining more knowledge or data [42, 30]. Techniques for the disentanglement of aleatoric and epistemic uncertainty are introduced by Valdenegro-Toro et al. [43], and Munikoti et al. [44] formulate mathematical expressions of these two kinds of uncertainty in the context of GNNs.
The identification of aleatoric and epistemic uncertainty must be built upon the context of the selected model and available data [45]. For example, flipping a coin is usually regarded as a purely random event, involving aleatoric uncertainty. However, suppose the entire physics process of flipping a coin is known, including the flipping force, the coin’s mass, volume, initial position, etc.. In that case, the process can be simulated, and whether the result would be a head or a tail is predictable. With such a physical simulation, randomness can be regarded as epistemic uncertainty and can be reduced or even eliminated.
To reduce the ambiguity mentioned above, we mathematically define the two sources of uncertainty and how they impact prediction reliability. Let $X\in\mathcal{X}\subseteq\mathbb{R}^{d}$ and $Y\in\mathcal{Y}\subseteq\mathbb{R}$ denote the input and output variables with realizations $x$ and $y$, respectively. The underlying data distribution is $P_{XY}$ and the data generation mechanism is $f_{0}$ such that $y=f_{0}(x)$.
@@ III-A Aleatoric Uncertainty
Formally, when the data-generating mechanism $f_{0}$ is conditioned on $X=x$, aleatoric uncertainty (AU) can be represented by the conditional distribution $P_{Y|X}(\cdot\mid x)$. The dispersion of this distribution quantifies the AU and is often measured by the conditional Variance $\mathrm{\text{Var}}_{Y|X}[Y\mid x]\in\mathbb{R}$. . Gruber et al. [29] identify four distinct sources of aleatoric uncertainty, categorized as follows.
Omitted Features. Omitted features are variables excluded during model training and prediction. They contribute to AU, since collecting more data without these missing features does not reduce the uncertainty they introduce. The omission is investigated in the context of causal models [46], confounder analysis [47], and model sensitivity analysis [48]. The omission may result from variables being unobservable or overlooked during data collection. Let $Z\in\mathcal{Z}$ denote an unobserved variable that influences the value of $Y$, the expectation of $Y|x$ is a weighted integration of $\mathbb{E}_{Y|X,Z}$ on $Z$:
@@
Gruber et al. [29] also point out that $\text{Var}_{Y|X,Z}[Y|x,z]\leq\text{Var}_{Y|X}[Y|x]$ with at least one $z\in\mathcal{Z}$, indicating that including the non-negligible variables will reduce AU.
Feature Errors. Feature errors represent inaccurate feature values in collected data, which are likely to occur with improper measurements, such as low-resolution cameras, low-quality audio recorders, and uncalibrated gauges [32]. These errors can cause biases and increased Variance in feature collection. Let $Z$ represent feature variables by faultless measurements and $X$ denote those by inaccurate measurements, assuming that $Y$ is independent of $X|Z$. The AU caused by feature errors is illustrated as
$\text{Var}_{Y|X}[Y|x]=$ $\mathbb{E}_{Z|X}\left[\text{Var}_{Y|Z}[Y|Z]|x\right]$ (8)
$\;+\text{Var}_{Z|X}\left[\mathbb{E}_{Y|Z}[Y|Z]|x\right],$
where the first term of the right-hand side approximates $\text{Var}_{Y|Z}(Y|Z)$ if $Y\perp\!\!\!\perp X|Z$, and the second term is an increment in AU caused by feature errors [29].
@@
This decomposition shows that the observed label distribution $P_{Y|X}$ can be viewed as a mixture over the true label distribution $P_{Z|X}$, modulated by the conditional noise model $P_{Y|X,Z}$, which captures the randomness introduced by the labeling process. In this case, uncertainty in $Y$ arises not only from inherent ambiguity in the input–output relationship, but also from label noise. As such, this contributes to aleatoric uncertainty, since it cannot be reduced merely by increasing the number of training samples without improving label quality.
Missing Data. Missing data indicates some entries in the dataset are incomplete [50]. This differs from omitted features, which refer to entire variables that are excluded from the dataset, rather than individual missing values within an included feature. Denoting $(x_{i},y_{i})$ the $i$-th sample of the dataset, $M_{i}$ is defined as the corresponding missing indicator of $(x_{i},y_{i})$. Specifically, $M_{i}=0$ means the sample is complete, and $M_{i}=1$ otherwise. With a bias factor ${\Pr(M=0|y,x)}/{\Pr(M=0|x)}$ [29], the influence from incompleteness can be expressed by
This equation reveals that missing data can introduce selection bias into the observed conditional distribution $P_{Y\mid X,M}$, which deviates from the true distribution $P_{Y\mid X}$. The uncertainty arising from this bias is aleatoric in nature, as it stems from incomplete observations that cannot be resolved simply by collecting more data unless the missingness mechanism is addressed or corrected.
III-B Epistemic Uncertainty
Epistemic uncertainty (EU) is from model selection and the lack of data and knowledge. EU can be categorized into model, approximation, and Out-of-Distribution (OOD) uncertainties [30].
Model Uncertainty. Model uncertainty refers to the difference between the ground truth mapping $f_{0}\in\mathcal{F}_{0}$ and the best possible predictor $f\in\mathcal{F}$. It is impossible to explore the entire function space $\mathcal{F}_{0}$, resulting in a gap between $f$ and $f_{0}$. Specifically, taking $l$ as the loss function during training, $f_{0}$ and $f$ can be explicitly defined as
[…]
Thereby, we can represent model uncertainty by $D(P_{XY},(\mathrm{id},f)_{\#}P_{X})$, where $D(\cdot,\cdot)$ is a user-specified discrepancy measure and $(\mathrm{id},f)_{\#}P_{X}$ is the pushforward measure of $P_{X}$ through the mapping $x\mapsto(x,f(x))$, i.e., the joint distribution of $(x,f(x))$ when $x\sim P_{X}$..
Approximation Uncertainty. Approximation uncertainty is defined as the uncertainty due to the limited number of $N$ training samples. In other words, $P_{XY}$, as well as $f$ calculated in Eq. (12) is not accessible in practice. Therefore, the induced predictor trained with finite samples is defined as
$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}l(y_{i},f(x_{i})),$ (13)
which allows us to quantify approximation uncertainty as $D(P_{XY},(\mathrm{id},\hat{f})_{\#}P_{X})$.
OOD Uncertainty. Unlike the previous two types, OOD (Out-of-Distribution) uncertainty arises at the inference stage when the deployed model encounters test data drawn from a distribution different from the training distribution [51, 52]. Formally, let $P_{XY}$ and $Q_{XY}$ denote the training and test distributions, respectively. OOD uncertainty occurs when $P_{XY}\neq Q_{XY}$.
This distributional shift can arise in real-world scenarios, such as changes in data collection environments (e.g., sensor drift in medical devices), evolving user behavior over time (e.g., preferences in recommender systems), or deployment in a new domain (e.g., applying a model trained on urban traffic data to rural areas). Since the model ${f}$ is trained to minimize the expected loss under $P_{XY}$ in Eq. (12), its performance and confidence estimates may be unreliable when evaluated under $Q_{XY}$. Therefore, OOD uncertainty can be formally quantified by $D(Q_{XY},(\mathrm{id},f)_{\#}Q_{X})$.
Model and approximation uncertainties are reducible by expanding the search space in the function space $\mathcal{F}$ and increasing the sample size $N$ [53, 54, 55]. In contrast, OOD uncertainty arises from distributional shifts between training and test data and cannot be mitigated by more data or a larger model alone. Instead, it requires robust training strategies, which will be introduced in Sec. V, to improve generalization under shifted conditions.